### v1

# Combating Partial Perception Deficit in Autonomous Driving with Multimodal LLM Commonsense

Abstract
Partial perception deficits can compromise autonomous vehicle safety by disrupting environmental understanding. Current protocols typically respond with immediate stops or minimal-risk maneuvers, worsening traffic flow and lacking flexibility for rare driving scenarios. In this paper, we propose LLM-RCO, a framework leveraging large language models to integrate human-like driving commonsense into autonomous systems facing perception deficits. LLM-RCO features four key modules: hazard inference, short-term motion planner, action condition verifier, and safety constraint generator. These modules interact with the dynamic driving environment, enabling proactive and context-aware control actions to override the original control policy of autonomous agents. To improve safety in such challenging conditions, we construct DriveLM-Deficit, a dataset of 53,895 video clips featuring deficits of safety-critical objects, complete with annotations for LLM-based hazard inference and motion planning fine-tuning. Extensive experiments in adverse driving conditions with the CARLA simulator demonstrate that systems equipped with LLM-RCO significantly improve driving performance, highlighting its potential for enhancing autonomous driving resilience against adverse perception deficits. Our results also show that LLMs fine-tuned with DriveLM-Deficit can enable more proactive movements instead of conservative stops in the context of perception deficits.
footnotetext: ${}^{\text{1}}$niversity at Buffalo, SUNY. ${}^{\text{2}}$niversity of Notre Dame. Correspondence to: Yuting Hu $<$yhu54@buffalo.edu$>$, Jinjun Xiong $<$jinjun@buffalo.edu$>$
Preprint.
1 Introduction
Autonomous driving has transformed the transportation landscape with its rapid evolution these years. Modern autonomous driving systems rely heavily on sensor data inputs from cameras and LiDARs, processed by deep neural networks, to enable perception, understanding, and interaction with the environment [17, 26, 7, 25]. Since perception directly influences critical driving decisions, partial information deficits caused by sensor failures or attacks can be fatal, leading to catastrophic consequences by impairing essential functions such as object detection and spatial sensing  [8, 31, 39, 46].
Figure 1: With perception deficits of safe-critical objects, the deep learning-based AV agents tend to take over-conservative or unsafe maneuvers. In contrast, LLM-RCO draws on LLM commonsense knowledge to infer possible hazards in the deficit area and plan appropriate maneuvers to solve the context.
Considering severe driving scenarios with perception deficits, as illustrated in Fig. 1, the loss of safety-critical object information can undermine driving safety. Conventional safety protocols often default to immediate stopping or minimal maneuvers, which can unnecessarily impede traffic flow and still fall short when confronting partial information deficits in rare driving scenarios. Moreover, most fail-safe strategies still depend on complete perception inputs, which can be unreliable under partial perception deficits, limiting their effectiveness [45]. Human drivers demonstrate that safe navigation is possible with limited visibility by leveraging commonsense reasoning and experience to distinguish or infer critical visual information [14]. For instance, obstructions in peripheral areas may not impact immediate driving decisions, and cautious proceeding remains viable when maintaining sufficient physical distance from deficit regions or when these deficit regions show temporal stability. This human-inspired insight suggests that autonomous systems should conduct proactive context-aware operations while ensure safety. While commonsense plays a crucial role in human judgment to ensure a smooth and safe driving experience, current autonomous driving research has yet to explore how to integrate commonsense into decision-making under such challenging conditions.
To address this challenge, we turn our attention to human-level intelligent multimodal large language models (LLMs), which have demonstrated remarkable reasoning capabilities and extensive commonsense knowledge across diverse applications [18, 51, 37]. Trained on vast web-scale corpora, multimodal LLMs inherently encode a wealth of traffic rules and driving knowledge, potentially equipping them with a strong driving commonsense. However, effectively leveraging this knowledge for autonomous driving remains an open question. It is not as simple as designing a prompt, particularly when driving conditions vary under different partial perception deficits. Moreover, unlike end-to-end LLM driving agents [41, 29], which can be trained through imitation learning from rule-based experts, there is no well-established expert to provide optimal strategies for vehicle movement under incomplete perception. Most fail-safe mechanisms simply default to an immediate stop, making it challenging to collect training data for an end-to-end deep learning model that directly guides autonomous vehicle movement in scenarios with perception deficits.
In this paper, we present LLM-Guided Resilient Control Override (LLM-RCO), a novel framework that harnesses the commonsense knowledge and reasoning capabilities of multimodal large language models to enhance the resilience and safety of autonomous vehicle control under partial perception deficits. Specifically, when camera vision is compromised, LLM-RCO overrides autonomous vehicle control to mitigate the risk of hazardous actions resulting from erroneous predictions on compromised perception data, ensuring safer operation in adverse conditions. Inspired by human driver’s chain of thought when facing perception deficits, our framework decomposes this reasoning process into four key LLM-enhanced modules that dynamically interact with the driving environment: (1) Hazard Inference Module, which assesses potential hazards within deficit regions with past camera frames; (2) Short-Term Motion Planner, which creates a variable-length sequence of action-condition pairs adapting to the driving context and hazard inference results; (3) Action Condition Verifier, which evaluates the consistency of information deficits and immediate hazards in the current observation to confirm action feasibility, enabling safe and reliable vehicle control actions; (4) Safety Constraints Generator, which generates vehicle control limits based on safety-critical factors such as weather, lighting conditions, and traffic density; Additionally, we construct the DriveLM-Deficit dataset based on DriveLM-GVQA [42] to fine-tune the LLM for hazard inference and planning, enhancing the subsequent short-term motion planning process. To our knowledge, this paper is the first to provide an LLM-based framework for autonomous driving against perception deficits.
Contributions. We make the following contributions:
  • We propose LLM-RCO, a framework that deconstructs the human driver’s chain of thought when facing perception deficits into four LLM-fostered modules, enabling proactive vehicle movement.
  • We construct DriveLM-Deficit, a dataset composed of 53, 895 driving videos with perception deficits and annotations of hazards and motion planning. We benchmark the performance of advanced LLMs on DriveLM-Deficit, highlighting their limitations.
  • We fine-tune Qwen2-VL-2B-Instruct on the DriveLM-Deficit dataset using LoRA to enhance hazard inference and motion planning of LLM-RCO.
  • We validate LLM-RCO’s effectiveness in closed-loop environments using the CARLA [12] simulator separately with TransFuser [34] and InterFuser [40] agents, demonstrating that LLM-RCO consistently enhances driving scores by adopting a cautious yet proactive driving style.
Figure 2: LLM-RCO framework overview.
2 Related Works
2.1 LLM in Autonomous Driving
Recent advancements have integrated large language models (LLMs) into autonomous driving systems, offering improvements across multiple areas. LLMs have been applied as autonomous driving agents in closed-loop scenarios, demonstrating their potential for independent task handling in controlled environments [29, 30, 19, 27, 41]. They also play a critical role in enhancing autonomous driving security by detecting and mitigating adversarial attacks, making vehicles more resilient to malicious disruptions [22, 43, 10, 3, 2]. Furthermore, LLMs contribute to explainability, providing insights into decision-making processes that foster safer and more transparent vehicle control operations [9, 33, 49]. In addition, embedding LLMs into existing autonomous driving technique stacks enables commonsense, knowledge-based decision-making, enhancing autonomous vehicle’s ability to make context-aware driving decisions [11, 38, 48, 13]. Despite these advancements, there remains a significant gap in addressing motion planning under conditions of partial perception deficit like attacks or sensor failures, where LLMs could strengthen autonomous driving resilience in real-world scenarios.
2.2 Fail-Safe Strategies of Autonomous Driving
Fail-safe strategies in autonomous driving rely on multi-sensor fusion [26, 6], safe degraded mode [28, 20], and emergency stop or pull-over maneuvers [5, 15] to ensure safety under perception failures and sensor malfunctions. However, these approaches have notable limitations: rule-based protocols struggle with long-tail scenarios, emergency stops disrupt traffic flow, and perception-dependent fail-safe motion planning remains vulnerable to sensor failures and adversarial attacks [28, 35, 23, 32]. While some AV frameworks use hard-coded safety control rules, such as overriding stops in low-confidence traffic light detections, these rigid mechanisms lack adaptability in real-world conditions [36, 50, 24]. A more robust approach should leverage context-aware decision-making rather than fixed rule-based overrides. This motivates LLM-RCO, which aims to provide a more effective and adaptable solution for handling perception failures in autonomous driving systems.
3 LLM-RCO
In this section, we propose the LLM-RCO, which features the incorporation of LLM commonsense in multiple steps of planning and reasoning processes to enhance the resilience and safety of autonomous vehicle control strategies.
3.1 Preliminary
We first define the final executed action of autonomous driving at time $t$ as $\mathcal{A}_{t}$. Action contains three vehicle control parameters: throttle, brake, and steering wheel, i.e., $\mathcal{A}_{t}=[\textit{throttle}_{t},\textit{brake}_{t},\textit{steer}_{t}]\in[0,1]\times[0,1]\times[-1,1]$. We then define an action sequence as a collection of pairs, each consisting of a tentative candidate action and its execution condition, spanning from time $t$ to time $t+n$, denoted as $\textit{S}_{t}=\{(\mathcal{C}_{t},\mathcal{A}_{t}),\cdots,(\mathcal{C}_{t+n},\mathcal{A}_{t+n})\}$. Finally, we define the environmental information at $t$ as $E_{t}=[\textit{Perception}_{t},\textit{Navi}_{t},\textit{Surrounding}_{t}]$, where $\textit{Perception}_{t}=\{I^{L}_{t},I^{F}_{t},I^{R}_{t}\}$ is the real-time images separately collected by left, front, and right camera sensors; $\textit{Navi}_{t}$ is the navigation contains the coordinate of the target point, current direction, and road geometry; and $\textit{Surrounding}_{t}$ is the surrounding environmental parameters that describe traffic, weather and daylight information. We define the safety conditions of the vehicle at time $t$ as $\textit{SC}_{t}$, describing constraints such as the maximum following distance and so on.
3.2 Framework Overview
Unlike the LLM-based autonomous driving agents that employ continuous, real-time action inference, our LLM-RCO framework adopts a proactive, plan-ahead strategy. By reducing the frequency of interactions with the LLM, it lowers inference latency, leading to more efficient and smoother driving. Fig. 2 illustrates overall the LLM-RCO framework, which consists of two phases: planning and acting.
(1) When planning, if $\textit{S}_{t}$ is empty, LLM-RCO updates $n$-steps in $\textit{S}_{t}$ based on the perception context and navigation:
$\textit{S}_{t}=F_{\text{plan}}(\textit{Navi}_{t},\textit{Percetion}_{\{t-k,\cdots,t\}}).$ (1)
LLM-RCO planning first launches a Hazard Inference Module (Section 4.1), an LLM agent that takes multiple frames of sensor context information $\textit{Percetion}_{\{t-k,\cdots,t\}}$, and then concludes the potential hazard and its movement into $\textit{Hazard}_{t}$ in Eq. (3). Then, a Short-term Motion Planning Module (Section 4.2), another LLM agent takes concluded hazard information $\textit{Hazard}_{t}$, navigation information $\textit{Navi}_{t}$ and current time perception information $\textit{Perception}_{t}$, and then generates a variable-length (we denote as $n$) condition-action sequence as $\textit{S}_{t}$.
(2) When acting, LLM-RCO reads condition-action pairs from $\textit{S}_{t}$ sequentially, referencing environment information:
$\begin{cases}(\mathcal{C}_{t},\mathcal{A}_{t})=F_{\text{act}}(\textit{S}_{t},E_{t})\\ \textit{S}_{t+1}=\textit{S}_{t}\setminus\{(\mathcal{C}_{t},\mathcal{A}_{t})\}.\end{cases}$ (2)
The condition-action pairs will first be verified by the Action Condition Verifier (Section 4.3), which determines whether the current environment perception $E_{t}$ is suitable for executing the condition-action $(\mathcal{C}_{t},\mathcal{A}_{t})$. If suitable, LLM-RCO executes the action $\mathcal{A}_{t}$ by the acting phase. If the Action Condition Verifier denies action $\mathcal{A}_{t}$ or the $\textit{S}_{t}$ is empty, then LLM-RCO will start a new round of planning.
4 LLM-RCO Key Components
4.1 Harzard Inference
Partial perception loss can severely impact driving when the AV system cannot detect potential hazards in compromised areas. Inferring these hazards in the perception deficit area enables the system to make more informed driving decisions. We define a hazard based on the object and its movement. Moreover, multiple hazards (denoted as $N_{t}$) may exist simultaneously. Formally:
$\textit{hazard}_{t}=\{[\textit{object}_{i},\textit{motion}_{i}]\arrowvert_{i=1}^{N_{t}}\}.$ (3)
In LLM-RCO, the hazard inference module is essential for ensuring safe driving decisions by compensating for perception loss. Errors in hazard inference can adversely affect short-term planning. To address this, we fine-tune LLMs to jointly infer potential hazards within compromised image regions and the corresponding short-term planning strategies mentioned in Section 4.2, as the pre-trained LLMs have already learned robust visual and linguistic features from large-scale datasets. Assuming that sensor failures or attack patterns remain consistent in the short term, tracking the temporal motion of the unaffected surroundings provides vital insights into their evolving dynamics, enabling the LLM to produce more accurate hazard predictions. Consequently, we input a sequence of the past $k$ multi-view images from the cameras, allowing the model to infer potential hazards and appropriate motion strategy in the deficit region based on historical context.
$(\textit{hazard}_{t},\textit{plan}_{t})=\text{LLM}_{\text{HI}}(\textit{Percetion}_{\{t-k,\cdots,t\}}).$ (4)
[…]
4.2 Short-term Motion Planing
Figure 3: We show two cases of movements using LLM-RCO under Move and Stop-Observe-Move strategies.
Short-term Motion Planner employs two distinct operational strategies $plan$ that emulate human-like adaptability: Move and Stop-Observe-Move. As demonstrated in Fig. 3, under Move strategy, the vehicle proceeds cautiously with short-term motions plan ahead, while Stop-Observe-Move requires temporary halting to gather additional information before proceeding to ensure safety in uncertain situations. Under the Move strategy, the system generates an adaptive-length sequence of planning steps. For Stop-Observe-Move strategy, when the LLM determines that an immediate stop is necessary, it specifies both the waiting duration $wait$ and move trigger conditions $\mathcal{C}_{\text{move}}$ to initiate a new round of motion planning for the vehicle’s movement. During the waiting period, only ‘stop’ actions are added into $\textit{S}_{t}$.
$\textit{S}_{t}=\text{LLM}_{\text{SMP}}([\textit{hazard}_{t},\textit{plan}_{t},\textit{Navi}_{t},\textit{Perception}_{t}]).$ (5)
Different from conventional defense strategies that only output driving control actions, LLM-RCO tasks the LLMs to generate an execution condition for each action to enhance driving safety and allow better adaptation to dynamic road conditions. In $\textit{S}_{t}$, each element from $(\mathcal{C}_{t},\mathcal{A}_{t})$ to $(\mathcal{C}_{t+n},\mathcal{A}_{t+n})$ is a condition-action pair, indicates that the action will be executed only when the condition is met. We define conditions as combinations of ‘consistent_deficit’ with either ‘no_immediate_hazard’ or ‘immediate_hazard’. When the observed deficits are inconsistent across frames, it implies extreme uncertainty about potential hazards arising from perception deficits. In such cases, the LLM is tasked with real-time motion inference by immediate replanning. Conversely, if deficit patterns remain consistent across frames, the LLM can utilize unaffected perception information and hazard inference results to generate proactive action sequences. As illustrated in Fig. 4, under consistent deficits, the LLM may plan more stable actions to enable smooth driving if there is no immediate hazard or opt for more aggressive maneuvers such as sudden brake in the presence of an immediate hazard.
Figure 4: Illustration of inconsistently/consistently deficits across frames in various hazard cases. Left: Scenarios of Inconsistent Deficits: (a) spatial shift, (b) quantity realignment, and (c) deficit disappearance. Right: Scenarios of Consistent Deficits: (a) no immediate hazard, (b) immediate hazard due to a large deficit size, (c) immediate hazard from nearby objects in unaffected regions.
4.3 Action Condition Verifier
At each timestep, LLM-RCO’s acting phase reads and executes the next action $\mathcal{A}_{t}$ in $\textit{S}_{t}$. To ensure the safe execution of plan-ahead actions, we design a rule-based action condition verifier to check $\mathcal{C}_{t}$ based on real-time perception observation. In LLM-RCO, we task LLM to generate a proactive action sequence to reduce LLM inference load. However, as the road conditions might be highly dynamic, we introduce a deficit consistency check to enable real-time action inference by motion replanning. As shown in Fig. 4, the deficit consistency is evaluated by examining the temporal quantity and spatial shifts of deficits observed across frames. We compare deficit regions across image frames in terms of both their quantity and relative spatial positions against the background. If the quantity of deficits does not align or if the spatial shift of deficit regions exceeds a predefined threshold, the condition check will fail, prompting instant motion replanning. If the consistency check passes, we assess immediate hazards by analyzing the ratio of deficit regions and detected traffic objects (e.g., cars, trucks, buses, bicycles, pedestrians, motorcycles using YOLOv11) to the camera image. This ratio, computed as the area of deficit regions and traffic object bounding boxes relative to the total image size, serves as a hazard proximity indicator. If it exceeds a predefined threshold (set to 0.05), it indicates an immediate hazard near the vehicle.
4.4 Safety Constraint Module
In LLM-RCO, the Safety Constraint Module adjusts the action $\mathcal{A}_{t}$ from the Action Condition Verifier to ensure compliance with safety constraints. As safety constraints depend on various factors, for example, the maximum following distance in snowy conditions differs from that in normal conditions, we leverage the extensive commonsense knowledge of LLMs to improve driving safety. Specifically, based on the driving context including weather, daylight, traffic density, road geometry, etc., we task LLMs to generate a set of loose safety constraints $\mathcal{SC}_{t}$ to enhance vehicle control safety. As driving conditions remain stable over short time periods, we don’t need real-time inference for $\mathcal{SC}_{t}$ generation. Due to space limitations, more implementation details are provided in the Appendix A.1.
Figure 5: Driving-Oriented Fine-tuning Scheme.
5 Driving-Oriented Fine-tuning
To deliver straightforward optimal driving actions to LLM, we propose fine-tuning an LLM for hazard inference and strategy selection to guide subsequent short-term motion planning. The fine-tuning process involves three tasks: (1) Object Inference, which identifies objects within deficit regions using contextual cues from unaffected parts of the image; (2) Motion Inference, which analyzes motion behaviors in deficit regions based on historical perception frames; and (3) Planning Strategy, which decides between “move” or “stop-observe-move” approaches using insights from the previous tasks. For example, the system may choose “move” with the cue “Bicycle, oncoming at a constant speed,” or “stop-observe-move” with the cue “Pedestrian, crossing to the other side of the road.” Since these three tasks depend on different aspects of the image information, as in Fig. 5, we use three Low Rank Adaptation (LoRA) [16] modules to adapt the model to these tasks.
Figure 6: DriveLM-Deficit Dataset.
We process GVQA dataset from DriveLM-Carla [42] for fine-tuning, which is collected using PDM-Lite [4] expert and achieves 100% completion with zero infractions. GVQA dataset comprises 1.6M QA pairs along with sensor data, keyframes, and frame-level QAs covering perception, prediction, and planning. It employs a rule-based annotation pipeline to determine the ego vehicle’s braking status using privileged information about objects, hazards, and scene measurements. Building on DriveLM-GVQA, we construct the DriveLM-Deficit dataset tailored to our specific case. DriveLM-Deficit consists of 53, 895 videos, each composed of five frames sampled every two steps, forming a 5-second clip at 1 fps. It includes deficits related to traffic lights, traffic signs, pedestrians, and bicycles for fine-tuning. The statistics of DriveLM-Deficit is shown in Fig. 6, and more details are in Appendix A.3.
6 Experiments
Fine-tuning Details. We choose to supervise fine-tune a lightweight pre-trained vision language model “Qwen2-VL-2B-Instruct” [47] with LoRA using DriveLM-Deficit. we only add LoRA to the “q_proj” and “v_proj” attention layers, and we set the rank to 4 for all experiments.
LLM-RCO Implementation. To control the autonomous agent, we define high-level actions with driving behavior (move forward, stop, change lane to left, change lane to right, turn left, turn right) and speed control (constant speed, deceleration, quick deceleration, deceleration to zero, acceleration, quick acceleration), the action tokens are mapped to control parameters in a deterministic manner based on the navigation to the next target point, more details are provided in Appendix A.1. We employ GPT-4o-mini [1] and fine-tuned Qiwen2-VL-2B as LLM-RCO backbone models.
Perception Deficit Scenarios. To assess the performance of autonomous driving systems under challenging conditions, we designed experiments that simulate different scenarios of perception deficit regarding safety-critical objects. Specifically, using YOLO11 [21], we track and mask safety-critical objects in real-time, including traffic lights, stop signs, pedestrians, and bicycles. When these objects are masked out, control transfers to LLM-RCO.
Carla Simulator. Our implementation employs CARLA (version 0.9.10.1) along with the Transfuser (TF) and Interfuser (IF) models, which integrate both image and LiDAR data for end-to-end vehicle control. Transfuser incorporates rule-based stuck detection and employs a force-move command within its control prediction. Similarly, Interfuser features a rule-based force-move mechanism integrated into its safe controller module. In our setup, both force-move mechanisms remain at their original default settings to enable a fair comparison with LLM-RCO. Experiments are conducted on the Longest6 benchmark, which comprises 36 routes averaging 1.5 km in length and featuring high densities of dynamic actors in diverse environmental conditions.
(a) Hazard inference.
(b) Motion strategy.
Figure 7: Accuracy of various LLM backbones on Drive-Deficit.
Metrics. We adopt three primary metrics employed by the CARLA Leaderboard for evaluations: Route Completion (RC), Infraction Score (IS), and Driving Score (DS). The IS typically accounts for collisions and traffic rule violations, including running red lights and stop signs. In the experiments, we modified the IS calculation correspondingly in the deficit scenarios of traffic lights and traffic signs by excluding red lights and stop sign violations from scoring. The DS is calculated as the product of RC and IS, reflecting both driving progress and safety. We report the metrics for three runs. Additionally, we report the Average Speed (AS) to assess driving efficiency, which is calculated as the total route length divided by the game time. Note that the game time is counted by the ticks in the simulation, and the model inference time is not factored into this count.
Table 1: Comparison of driving performance with unlimited game time. denotes the highest driving score.
Experiment Metric Traffic Light Stop Sign Pedestrian Bicycle
TF IF TF IF TF IF TF IF
NO Visual Deficit RC $100$ $100$ $100$ $100$ $100$ $100$ $100$ $100$
IS $0.13\pm 0.06$ $0.71\pm 0.08$ $0.13\pm 0.06$ $0.71\pm 0.08$ $0.13\pm 0.06$ $0.71\pm 0.08$ $0.13\pm 0.06$ $0.71\pm 0.08$
DS $13.18\pm 6.17$ $71.30\pm 8.09$ $13.18\pm 6.17$ $71.30\pm 8.09$ $13.18\pm 6.17$ $71.30\pm 8.09$ $13.18\pm 6.17$ $71.30\pm 8.09$
AS $1.55\pm 0.28$ $2.14\pm 0.07$ $1.55\pm 0.28$ $2.14\pm 0.07$ $1.55\pm 0.28$ $2.14\pm 0.07$ $1.55\pm 0.28$ $2.14\pm 0.07$
Visual Deficit LLM-RCO RC 100 $94.30\pm 5.01$ 100 $95.39\pm 5.26$ 100 $94.99\pm 5.01$ 100 100
IS $0.05\pm 0.04$ $0.29\pm 0.05$ $0.02\pm 0.01$ $0.23\pm 0.07$ $0.003\pm 0.001$ $0.16\pm 0.08$ $0.07\pm 0.02$ $0.36\pm 0.07$
DS $4.88\pm 3.76$ $27.44\pm 3.43$ $2.18\pm 0.97$ $21.20\pm 5.67$ $0.32\pm 0.11$ $15.12\pm 6.48$ $6.54\pm 2.10$ $36.47\pm 6.93$
AS $1.76\pm 0.07$ $1.78\pm 0.01$ $1.87\pm 0.02$ $2.44\pm 0.23$ $1.95\pm 0.21$ $1.05\pm 0.13$ $1.88\pm 0.07$ $2.27\pm 0.06$
Visual Deficit LLM-RCO HI-GPT,STP-GPT RC $94.53\pm 0.24$ $98.31\pm 4.28$ $100$ $100$ $100$ $94.10\pm 3.45$ $100$ $95.41\pm 0.11$
IS $0.11\pm 0.05$ $0.32\pm 0.06$ $0.03\pm 0.01$ $0.22\pm 0.07$ $0.019\pm 0.004$ $0.21\pm 0.04$ $0.09\pm 0.03$ $0.43\pm 0.07$
DS $10.97\pm 3.96$ 30.70$\pm$5.53 2.67$\pm$0.88 $21.19\pm 6.84$ $1.87\pm 0.37$ 19.99$\pm$4.76 $9.07\pm 2.71$ 41.80$\pm$6.95
AS $1.97\pm 0.04$ $1.89\pm 0.04$ $1.94\pm 0.02$ $1.72\pm 0.03$ $1.48\pm 0.03$ $0.94\pm 0.17$ $1.40\pm 0.07$ $0.86\pm 0.01$
Visual Deficit LLM-RCO HI-Qiwen,STP-GPT RC $94.15\pm 0.33$ $99.10\pm 5.15$ $100$ $100$ $100$ $95.30\pm 4.17$ $100$ $100$
IS 0.12$\pm$0.04 $0.31\pm 0.06$ 0.03$\pm$0.01 $0.23\pm 0.07$ 0.023$\pm$0.006 $0.19\pm 0.03$ $0.09\pm 0.02$ $0.42\pm 0.08$
DS 11.18$\pm$3.87 $30.4\pm 5.66$ 2.59$\pm$0.73 22.65$\pm$6.91 2.34$\pm$0.64 18.71$\pm$4.05 9.13$\pm$2.45 $41.54\pm 8.10$
AS $2.08\pm 0.03$ $1.82\pm 0.06$ $1.96\pm 0.03$ $1.73\pm 0.02$ $1.15\pm 0.02$ $0.92\pm 0.33$ $1.42\pm 0.08$ $0.97\pm 0.03$
Table 2: Comparison of driving performance with a 30-minute system time limit. denotes the highest driving score.
Experiment Metric Traffic Light Stop Sign Pedestrian Bicycle
TF IF TF IF TF IF TF IF
NO Visual Deficit RC $97.35\pm 3.89$ $99.13\pm 1.41$ $97.35\pm 3.89$ $99.13\pm 1.41$ $97.35\pm 3.89$ $99.13\pm 1.41$ $97.35\pm 3.89$ $99.13\pm 1.41$
IS $0.14\pm 0.06$ $0.71\pm 0.03$ $0.14\pm 0.06$ $0.71\pm 0.03$ $0.14\pm 0.06$ $0.71\pm 0.03$ $0.14\pm 0.06$ $0.71\pm 0.03$
DS $13.52\pm 4.93$ $70.62\pm 5.15$ $13.52\pm 4.93$ $70.62\pm 5.15$ $13.52\pm 4.93$ $70.62\pm 5.15$ $13.52\pm 4.93$ $70.62\pm 5.15$
Visual Deficit LLM-RCO RC $30.19\pm 5.32$ $44.90\pm 0.55$ $37.03\pm 0.18$ 45.33$\pm$0.21 37.11$\pm$0.06 $42.20\pm 1.07$ 38.22$\pm$0.04 $44.70\pm 0.93$
IS 0.6$\pm$0.11 0.63$\pm$0.09 $0.14\pm 0.03$ $0.68\pm 0.10$ $0.20\pm 0.05$ $0.36\pm 0.09$ $0.17\pm 0.08$ $0.70\pm 0.06$
DS $17.89\pm 3.53$ $31.43\pm 3.07$ $5.0\pm 1.16$ $32.73\pm 3.89$ $7.48\pm 1.87$ $16.12\pm 3.82$ $6.43\pm 3.11$ $31.29\pm 2.82$
Visual Deficit LLM-RCO HI-GPT,STP-GPT RC 58.57$\pm$3.07 $56.11\pm 1.92$ 37.17$\pm$0.45 $43.15\pm 1.22$ $36.20\pm 0.16$ $41.10\pm 1.32$ $32.50\pm 0.42$ $42.92\pm 0.91$
IS 0.32$\pm$0.06 $0.60\pm 0.07$ $0.24\pm 0.03$ 0.78$\pm$0.09 0.28$\pm$0.02 0.42$\pm$0.09 $0.80\pm 0.13$ 0.78$\pm$0.07
DS 18.58$\pm$3.51 $33.43\pm 4.18$ $8.95\pm 0.72$ 33.70$\pm$4.09 9.94$\pm$0.81 18.79$\pm$4.66 26.26$\pm$4. $32.88\pm 3.01$
Visual Deficit LLM-RCO HI-Qiwen,STP-GPT RC 61.02$\pm$4.88 58.44$\pm$2.31 $36.39\pm 0.87$ $43.60\pm 1.27$ 40.12$\pm$0.70 44.81$\pm$1.33 $30.52\pm 0.62$ 46.80$\pm$1.21
IS $0.28\pm 0.04$ $0.58\pm 0.04$ 0.29$\pm$0.03 $0.74\pm 0.08$ $0.28\pm 0.03$ $0.41\pm 0.08$ 0.81$\pm$0.07 $0.76\pm 0.06$
DS $17.13\pm 2.44$ 33.90$\pm$2.34 10.60$\pm$1.17 $32.26\pm 3.51$ 11.03$\pm$1.66 $18.27\pm 4.40$ $25.02\pm 2.94$ 35.60$\pm$4.56
7 Results
7.1 Fine-tuning results on DriveLM-Deficit
In Fig. 7, we evaluate the hazard inference and motion planning accuracy of GPT-4o-mini, Gemini-2-flash [44], Qiwen2-VL, and fine-tuned Qiwen2-VL on the Drive-Deficit dataset. As shown in Fig.7(a), even advanced LLMs struggle to infer possible objects in occluded or missing image regions under driving scenarios. While they excel at identifying traffic lights and stop signs, they perform poorly in rare cases involving pedestrians and bicycles. To address this limitation, we fine-tuned the open-source Qiwen2-VL model for hazard inference and it achieves higher accuracy on both pedestrians and bicycles than GPT-4o-mini and Gemini2-flash. In Fig. 7(b), we evaluate the LLMs’ ability to infer appropriate motion plan based on past camera frames under perception deficits. The results indicate that advanced LLMs often opt to conservative stop even in scenarios where movement is feasible, which is reflected in lower prediction accuracy for “move” motion strategy across all deficit cases.
Figure 8: LLM-RCO performance under different action step length limit. Higher $\Delta RC$, $\Delta IS$, and $\Delta DS$ indicate better performance.
Figure 9: Speed control under various perception deficit scenarios.
7.2 System Results
Overall Results. We report the evaluation results of the LLM-RCO against various perception deficits in two experimental settings: without game time limit shown in Table 1 and with 30-min system time limit shown in Table 2. The game time excludes model inference time, and system time accounts for inference time. In Table 1, we report the average speed (AS) to evaluate the stop rate in the AV movement, where a lower AS indicates more frequent stops during travel. The results in both settings reveal that perception deficits significantly impair autonomous driving performance, leading to remarkably lower driving scores than scenarios without deficits. For Transfuser, the average speed steadily increases with perception deficits, indicating that the loss of critical hazard information leads to inattentive driving. In contrast, for the Interfuser, we observe that the loss of traffic light information results in overly conservative driving due to incorrect traffic light state predictions, causing it to remain stationary until forcibly moved by its rule-based safe controller.
When the LLM-RCO framework intervenes in vehicle control of these adverse scenarios, as shown in Table 1, we observe that the driving score shows improvement with a little compromised route completion score with unlimited game time. In stop sign, pedestrian, and motorcycle cases, the ego vehicle makes more stops, reflected by a decrease in average speed. This cautious driving control of LLM-RCO leads to a higher driving score. For traffic light scenarios, LLM-RCO increases the average speed by reducing unnecessary stops, addressing the issue where incorrect traffic light predictions in Interfuser cause the vehicle to remain stuck until forcibly moved. Moreover, considering the inference time cost of LLM, as shown in Table 2, despite the lower infraction score of LLM-RCO in the traffic light deficit scenario, where the longer completed routes increase exposure to potential risks for transfuser, the infraction and driving scores of LLM-RCO present a consistent increase in all other perception deficit scenarios. The results also demonstrate the fine-tuned Qiwen with more accurate hazard inference and motion planning can reduce stops when the game time is unlimited to enable a complete route and trigger more route completion with limited system time.
Speed Control Analysis. In Fig. 9, we analyze the speed control decisions made by LLM-RCO when overriding the autonomous agent in various perception deficit scenarios. The results are reported with Transfuser on route0 of Town01 in longest6. For scenarios with visual deficits of critical objects, the percentage of different types of speed control remains relatively consistent across test scenarios: approximately $50\%$ of actions maintain a constant speed, around $40\%$ involve deceleration, and only about $10\%$ involve acceleration, indicating that LLM-RCO adopts a cautious driving style, prioritizing safety. Additionally, we observe that the percentage of acceleration actions is lowest in the pedestrian masking scenario, in accordance with rigorous human safety.
Short-term Motion Planning Analysis. In Fig. 8, we examine the performance of LLM-RCO with different limits on planning step lengths. Within the LLM-RCO framework, a maximum length limit is applied to short-term motion planning, enabling the LLM to reason through a variable number of actions based on the current driving context. We report the evaluation metric change with and without LLM-RCO intervention in Transfuser. Our results reveal that extending planning steps increases accident exposure and lowers infraction scores. While a moderate increase in the step limit can improve driving scores, infraction scores drop sharply at higher limits, indicating that excessive planning steps ultimately degrade performance. We also observe that LLM-RCO exhibits different control strategies for static objects (Traffic Lights, Stop Signs) versus moving objects (Bicycles, Pedestrians). When facing static objects, LLM-RCO tends to execute more proactive maneuvers, reflected in higher average speeds and greater route completion (see Tables 2 and 3). Moderately enlarging the planning step limit can further boost driving scores by improving route completion. In contrast, for moving objects, LLM-RCO typically makes conservative stops, which lowers average speeds and hinders route completion. Slightly extending the step length limit prompts the agent to make additional stops, thereby increasing driving scores with increasing infraction scores.
Abalation Study. In Table 3, we conduct ablation studies over the LLM-RCO components based on Transfuser and Interfuser: the Safety Constraint Generator (SC), Hazard Inference Module (HI), Short-Term Motion Planner (ST), and Action Condition Verifier (AC). The results demonstrate the effectiveness of each module. We report the ablation results in the scenario of motorcycle deficit. The results illustrate the contribution of each module to overall driving performance and safety. The test demonstrates that full configuration with all modules active achieves the highest infraction and driving scores. Especially, the core modules of LLM-RCO (HI, ST, AV) are especially necessary to enable valid control override.
Table 3: Ablation study on the LLM-RCO module design.
LLM-RCO TF IF
HI ST AV SC RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$
37.14 0.13 5.01 44.70 0.70 31.29
36.35 0.25 9.16 41.27 0.70 28.89
36.35 0.42 15.27 41.19 0.75 30.89
34.36 0.77 26.50 43.27 0.77 33.31
33.77 0.80 27.08 42.92 0.78 32.88
8 Conclusions
Focusing on a realistic scenario in which autonomous vehicles need to ensure safe driving and avoid congestion when their perception is partially compromised, in this paper, we present LLM-RCO, a novel vehicle control framework designed to address this challenge. Our framework harnesses the advanced reasoning capabilities and comprehensive commonsense knowledge inherent in multimodal Large Language Models to enhance autonomous vehicle resilience and safety. By inferring potential hazards within the deficit region and generating corresponding planning-ahead vehicle actions following human-like motion strategies, LLM-RCO can override the vehicle control of autonomous driving agents. Experimental validation demonstrates LLM-RCO’s performance to enable proactive movement despite visual data loss, underlining its potential for modern autonomous driving systems.
Supplementary Material
Appendix A Implementation Details
A.1 Autonomous Vehicle Control
The control parameters in the CARLA simulator include steering, throttle, and brake. We define action as a combination of driving behavior (e.g., move forward, stop, change lane to the left, change lane to the right, turn left, or turn right) and speed control (e.g., constant speed, deceleration, quick deceleration, deceleration to zero, acceleration, or quick acceleration). To translate speed control tokens into CARLA control parameters, we apply the mapping outlined in Table 4. For driving behaviors that influence the steering angle, we use CARLA’s PIDController of Autopilot to determine the exact steer parameter based on the ego vehicle’s relative position to the next navigation target point. If the LLM-generated action involves a direction change instruction in driving behavior and it aligns with the moving direction on the navigation map, the steering parameters are subsequently computed by the autopilot.
Speed Control Carla Control Parameter
constant speed $throttle_{t}=0.7,brake_{t}=0$
deceleration $throttle_{t}=max(0,throttle_{t-1}-0.2),brake_{t}=0.2$
quick deceleration $throttle_{t}=max(0,throttle_{t-1}-0.4),brake_{t}=0.4$
deceleration to zero $throttle_{t}=0,brake_{t}=0.8$
acceleration $throttle_{t}=min(1,throttle_{t-1}+0.2),brake_{t}=0$
quick acceleration $throttle_{t}=min(1,throttle_{t-1}+0.4),brake_{t}=0$
Table 4: Mapping of speed control commands to vehicle control.
A.2 Safety-Constraints Generation
The Safety Constraint Module adjusts the action $\mathcal{A}_{t}$ transported from the Action Condition Verifier to meet the safety constraints in conjunction with measurements of the current vehicle’s driving conditions. To accomplish this, the Safety Constraint Module obtains the precise measurements of the vehicle’s speed $v$, longitudinal acceleration $a_{\text{x}}$, yaw rate $\omega_{\text{z}}$, and distance from the vehicle in front $d_{\text{follow}}$ through built-in Inertial Measurement Unit and Speedometer. If these measurements satisfy the Trigger Condition in Table 5, we transform the action accordingly so that our vehicle meets the appropriate safety constraints. We utilize LLM to gauge the current safety constraints $\textit{SC}_{t}=\{v_{\text{max}},d_{\text{min}},ac_{\text{max}},de_{\text{max}},\psi_{\text{max}},d_{\text{brake}}\}$ as in Table 5 in conjunction with navigation information $\textit{Navi}_{t}$, surrounding information $\textit{Surrounding}_{t}$ and LiDAR sensor information $I_{t}^{\text{LiDAR}}$:
$\textit{SC}_{t}=\text{LLM}_{\text{SC}}([\textit{Navi}_{t},\textit{Surrounding}_{t},I_{t}^{\text{LiDAR}}])$ (6)
Safety Constraint Trigger Condition Action Transformation
Max Speed Limit $v\geq v_{\text{max}}$ $\textit{throttle}_{t}-\Delta_{\text{throttle}}$
Min Following Distance $d_{\text{follow}}<d_{\text{min}}$ $\textit{throttle}_{t}-\Delta_{\text{throttle}}$
Max Acceleration Limit $a_{\text{x}}>ac_{\text{max}}$ $\textit{throttle}_{t}-\Delta_{\text{throttle}}\cdot(a_{\text{x}}-ac_{\text{max}})$
Max Deceleration Limit $a_{\text{x}}<-de_{\text{max}}$ $\textit{brake}_{t}-\Delta_{\text{brake}}\cdot(-de_{\text{max}}-a_{\text{x}})$
Max Yaw Rate Limit $|\omega_{\text{z}}|>\psi_{\text{max}}$ $\textit{steer}_{t}\times\frac{\psi_{\text{max}}}{|\omega_{\text{z}}|}$
Min Braking Distance $\frac{v^{2}}{2\times de_{\text{max}}}>d_{\text{brake}}$ $\textit{brake}_{t}+\Delta_{\text{brake}}$
Table 5: Safety constraints and corresponding transformation of autonomous vehicle control parameters.
For each safety constraint, when the trigger condition is satisfied, we perform the corresponding transformation on the corresponding action $\mathcal{A}_{t}$’s manifold ($\textit{throttle}_{t}$,$\textit{brake}_{t}$ or $\textit{steer}_{t}$). We set two hyperparameters $\Delta_{\text{throttle}}$ and $\Delta_{\text{brake}}$ here, as the transformation factors for the throttle and the brake specifically. Thus, we get the final action manifold:
$\begin{split}\hat{\textit{throttle}}_{t}=&\textit{throttle}_{t}-\mathbf{1}_{v\geq v_{\text{max}}}\cdot\Delta_{\text{throttle}}\\ &-\mathbf{1}_{d_{\text{follow}}<d_{\text{min}}}\cdot\Delta_{\text{throttle}}\\ &-\mathbf{1}_{a_{\text{x}}>ac_{\text{max}}}\cdot\Delta_{\text{throttle}}\cdot(a_{\text{x}}-ac_{\text{max}}),\end{split}$ (7)
$\begin{split}\hat{\textit{brake}}_{t}=&\textit{brake}_{t}+\mathbf{1}_{\frac{v^{2}}{2\times de_{\text{max}}}>d_{\text{brake}}}\cdot\Delta_{\text{brake}}\\ &+\mathbf{1}_{a_{\text{x}}<-de_{\text{max}}}\cdot\Delta_{\text{brake}}\cdot(a_{\text{x}}-de_{\text{max}}),\end{split}$ (8)
$\hat{\textit{steer}}_{t}=\textit{steer}_{t}\times[1-\mathbf{1}_{|\omega_{\text{z}}|>\psi_{\text{max}}}\cdot(1-\frac{\psi_{\text{max}}}{|\omega_{\text{z}}|})]$ (9)
Compositing Eq. (7), (8), and (9), we arrive at the final action $\mathcal{A}_{t}$ that is executed at time $t$:
$\mathcal{A}_{t}=[\hat{\textit{throttle}}_{t},\hat{\textit{brake}}_{t},\hat{\textit{steer}}_{t}]$ (10)
Figure 10: Examples of brake in DriveLM-GVQA and ‘Stop-Observe-Move’ in DriveLM-Deficit.
Figure 11: Examples of ‘Move’ and ‘Stop-Observe-Move’ motion strategies when the front bicycle is not visible in DriveLM-Deficit.
A.3 DriveLM-Deficit Dataset
To fine-tune an LLM for hazard inference and motion planning under perception deficits, we post-process DriveLM-GVQA to create DriveLM-Deficit. Specifically, we construct videos using 5 consecutive camera frames from DriveLM-GVQA. The hazard detection and braking annotations of DriveLM-GVQA at the final frame are used to derive the hazard motion and planning strategy label, either “move” or “stop-observe-move”. If the last frame indicates braking, we assign the “stop-observe-move” label; otherwise, we set it to “move”. As shown in Fig.10 (a), in DriveLM-GVQA, the detected hazards are used to determine the brake status for question “Does the ego vehicle need to brake? Why?” For example, “the ego vehicle should stop because of the traffic light that is red.” or “The ego vehicle should stop because of the pedestrian that is crossing the road.” Otherwise, the answer is: ”There is no reason for the ego vehicle to brake.” which corresponds to “move” strategy in our case.
To construct DriveLM-Deficit, we utilize YOLOv11 to track and occlude critical objects in the video, considering two occlusion scenarios. First, as shown in Fig. 10 (b), we directly occlude labeled hazards from DriveLM-GVQA, such as a “stop sign.” In this case, the rationale for the “stop-observe-move” label is: ”The ego vehicle should stop because the invisible region may contain a stop sign.” Second, as shown in Fig. 10 (c), we occlude other objects, such as a “pedestrian” near a stop sign. In this scenario, the ego vehicle must still stop due to the stop sign, regardless of any hazards inferred in the occluded region. We show two data examples of “move” or “stop-observe-move” planning strategy with occluded bicycle in the perception in Fig. 11.
A.4 Prompt and Output Examples
Figure 12: Hazard inference example.
Figure 13: Short-term motion planning example.